A rectangular iron pipe has an external dimention 12cm and 10cm with the iron being 1cm thick. The mass of 1cm^3 of cast iron is 7.2g. Calculate the mass of 2-meter length of the pipe.
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52848)

.

Now the problem formulation is IDEAL.


Solution

The mass of the pipe is the product of its volume and the metal density.


The volume of the metal part is equal to the volume of the rectangular prism of the "height" of 2 meters 
with the cross section dimensions 12 and 10 centimeters MINUS the volume of the interior empty part, which is the rectangular prism 
of the same "height" of 2 meters and with the cross section dimensions 12 - 2*1 = 10 and 10-2*1 = 8 centimeters.


So, the volume of the metal is


    V = 200*12*10 - 20*10*8 = 22400 cubic centimeters.


The mass of the iron is  7.2*22400 = 161280 grams = 161.280 kilograms.    ANSWER.

Solved.



/\/\/\/\/\/\/\/

Math problem,  correctly formulated and presented,  is and should be a piece of art.

Every word must be on its right place and correctly used.  All the words should be correctly placed and connected each with others.
No ambiguity is allowed.  No discontinuities or gaps in logic line are allowed.  No excessive words are allowed.

Formulation of Math problem is an art and a science.

Learn using good examples from good sources . . .


One very wise scientist and teacher said once that problems for students are the same and play the same role as the fair tales for children
- via such problems the students learn the world around.



Answer by Alan3354(69443)

IDK anyone who would not assume the cross-section is circular when you say "pipe."
--------
Post it like this:
A rectangular pipe has a cross-section of 10 by 12 cms outside, and it's 1 cm thick.
etc.
---------
Makes it simple and understandable.
======================================
Pipe ---> a round object.
You could have said "a rectangular wave guide" with dimensions. tho they are copper, not iron.